# Streamlined Supergravity

### v1

@@ 2 Construction of streamlined supergravity with a nilpotent superfield
Streamlining, by definition, means making something more efficient and simple by removing unnecessary parts. Note that the nilpotent sector acts behind the scenes; our only goal is to find a simple and efficient way to produce the most general scalar potentials and Kähler potentials in the physical sector. In our case, it involves two simplifications in the nilpotent sector of the theory.
@@
Here the constant $W_{0}$ defines the gravitino mass, whereas the constant $F_{s}$ is a non-vanishing value of the $D_{s}W|_{s=0}=F_{s}$. We keep here a simple Kähler potential (5) and a simple superpotential (6), but our Kähler potential of physical scalars $K(z^{i},\bar{z}^{\bar{i}})$ is arbitrary. This choice of the simple $s$-sector is convenient and sufficient to construct supergravity with any desired potentials $V(z^{i},\bar{z}^{\bar{i}})$. Our choice of (6) simplifies the general expression for $F_{s}=D_{s}W$ used in Kallosh:2015sea and makes it possible to use a constant $F_{s}$ instead of a function $F_{s}(z^{i})$.
@@
Here $\dots$ include various fermions. We will present more details in the Appendix to describe the complete supergravity action, including scalars, fermions, and vectors. In supergravity in Kallosh:2015sea and here in eq. (2) in absence of fermions
@@ 3 Consistency issue of streamlined supergravity and its resolution
$K(T,\bar{T},s,\bar{s})=-3\alpha\ln(T+\bar{T})+\frac{F_{s}\bar{F}_{\bar{s}}}{(T+\bar{T})^{3\alpha}[V(T,\bar{T})]+3|W_{0}|^{2}(1-\alpha)}\,s\,\bar{s}\ ,$ (14)
@@
${{\cal L}(T,\bar{T})\over\sqrt{-g}}={R\over 2}-{3\alpha\over 4}\,{\partial T\partial\bar{T}\over({\rm Re}\,T)^{2}}-[V(T,\bar{T})]\ .$ (15)
@@
However, we can use local supersymmetry to gauge fix the fermion from the nilpotent multiplet to vanish,
@@
Therefore, the kinetic terms of the boson $s$ and the fermion $\psi^{s}$ with a negative $K_{s\bar{s}}$ are not harmful. There is no vacuum instability due to these fermions since their kinetic term is either absent in $\psi^{s}=0$ gauge, or involves terms with negative $K_{s\bar{s}}$ which are not quadratic in physical fermions since they contain coupling to moduli $z^{i}$. Thus the new supergravity defined by $K_{s\bar{s}}$ in eqs. (11) is consistent even in the models where $K_{s\bar{s}}<0$ at some points in moduli space.
@@ 4.1.1 E-models
For models with an exponential approach to the plateau, consider the case of E-models of $\alpha$-attractors Ferrara:2013rsa ; Kallosh:2013yoa . In streamlined supergravity, it is described by complex half-plane variables $T=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$ in hyperbolic geometry with the nilpotent superfield $S$ with $S^{2}=0$. The simplest models of this class have potentials with a plateau, which is approached exponentially fast at large $\varphi$. In particular, one may consider models (14), (15) with
$V_{E}(T,\bar{T})=\Lambda+V_{0}\Big(1-{T+\bar{T}\over 2}\Big)^{2n}=\Lambda+V_{0}\Big(1-e^{-\sqrt{2\over 3\alpha}\varphi}\Big)^{2n}\ .$ (24)
Here $T=\bar{T}=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$, and inflation occurs at large values of the canonically normalized field $\varphi$. Note that this potential does not depend on the axion field $a$. To stabilize the axion field at $a=0$, one may simply add to this potential a term $m^{2}(T-\bar{T})^{2}$.
These models have a potential with a plateau, which is approached exponentially fast at large $\varphi$. In particular, for $n=1$
$V(T,\bar{T})=V_{0}\Big(1-{T+\bar{T}\over 2}\Big)^{2}=V_{0}\Big(1-2e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)\ .$ (25)
Note that the factor 2 in the last equation can be absorbed by a redefinition (a shift) of the field $\varphi$. This model can also be interpreted as pole inflation Galante:2014ifa , with the kinetic terms having a pole at $T=\bar{T}\to 0$.
@@ 4.1.2 T-models
The new supergravity version of inflationary models in disk variables, based on the general equation (4), corresponds to
$K(Z,\bar{Z},s,\bar{s})=$ $-3\alpha\ln(1-Z\bar{Z})+\frac{|f_{s}|^{2}}{(1-Z\bar{Z})^{3\alpha}\,V(Z,\bar{Z})+3|W_{0}|^{2}(1-\alpha Z\bar{Z})}s\bar{s}\ .$ (28)
@@
The potential in disk variables with the stabilized axion and $Z=\bar{Z}=\tanh{\varphi\over\sqrt{6\alpha}}$ during inflation at large $\varphi$ is
$V(Z,\bar{Z})\to V_{0}(1-{1+Z\over 1-Z}+\dots)=V_{0}\Big(1-e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)$ (32)
This is a case of the pole inflation in disk variables, where the kinetic terms have a pole of the order $p=2$ at the boundary $Z=\bar{Z}\to-1$. We find the same exponential approach to the plateau in terms of a canonical field $\varphi$ as in the E-model.
The supergravity version of polynomial $\alpha$-attractors Kallosh:2022feu is also described in (14), (15), with the kinetic term with the pole of order $p=2$. The potential in this case has particular features near the pole, where, e. g. for $k=2,4$
$V_{\rm pole}^{\rm ln}(T,\bar{T})\to V_{0}\Big(1-{1\over\ln^{k}T}+\dots\Big)=V_{0}\Big(1-\Big({\sqrt{2\over 3\alpha}\varphi}\Big)^{-k}+\dots\Big)$ (33)
In this case, the plateau at large $\varphi$ is approached polynomially Kallosh:2022feu ; Kallosh:2025ijd .
4.2 Polynomial approach to the plateau and pole inflation with $p\neq 2$
Now we will describe pole inflation with kinetic terms that have a pole of order $p>2$. To do it, we will take the following function $K(T,\bar{T},s,\bar{s})$ defined in the general equation (4):
$K(T,\bar{T},s,\bar{s})={\mu^{2}\over 2}({T}\overline{T})^{2-p\over 2}+{f_{s}^{2}\over{V}\left({T},\overline{{T}}\right)e^{-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}}+W_{0}^{2}\left(3-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}\right)}\,s\bar{s}$ (34)
@@ 4.2 Polynomial approach to the plateau and pole inflation with $p\neq 2$
This is the case of the pole inflation Galante:2014ifa with $p>2$, where in the canonical variables
$V|_{\varphi\to\infty}\to V_{0}(1-\Big({\mu\over\varphi}\Big)^{k}+\dots)\,\quad k={2\over p-2}>0\ .$ (37)
@@ 5 Discussion
In this paper we presented an efficient and relatively simple version of the streamlined $N=1$ supergravity with chiral multiplets and a nilpotent multiplet $S^{2}=0$ with the general type Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ leading to any desired scalar potential $V(z^{i},\bar{z}^{\bar{i}})$, once the proper choice of the Kähler metric of the nilpotent multiplet is made. The total Kähler potential, superpotential, and the bosonic action of this new supergravity for scalars are given in eqs. (12), (13). We gave some examples of applications of our result relevant to cosmology.
@@ Acknowledgement
We acknowledge many stimulating discussions over the years on supergravity with nilpotent multiplets with J.J. Carrasco, S. Ferrara, D. Freedman, E. McDonough, D. Roest, M. Scalisi, A. Van Proeyen, T. Wrase, and Y. Yamada. This work was supported by the Leinweber Institute for Theoretical Physics at Stanford and by NSF Grant PHY-2310429.

### v2

@@ 2 Construction of streamlined supergravity with a nilpotent superfield
Streamlining, by definition, means making something more efficient and simple by removing unnecessary parts. Note that the nilpotent sector acts behind the scenes; our only goal is to find a simple and efficient way to construct the most general scalar potentials and Kähler potentials in the physical sector. In our case, it involves two simplifications in the nilpotent sector of the theory.
@@
The constant $W_{0}$ defines the gravitino mass, whereas the constant $F_{s}$ is a non-vanishing value of the $D_{s}W|_{s=0}=F_{s}$. We keep here a simple Kähler potential (5) and a simple superpotential (6), but our Kähler potential of physical scalars $K(z^{i},\bar{z}^{\bar{i}})$ is arbitrary. This choice of the simple $s$-sector is convenient and sufficient to construct supergravity with any desired potentials $V(z^{i},\bar{z}^{\bar{i}})$. Our choice of (6) simplifies the general expression for $F_{s}=D_{s}W$ used in Kallosh:2015sea and makes it possible to use a constant $F_{s}$ instead of a function $F_{s}(z^{i})$.
@@
Here $\dots$ denote various terms containing fermions. We will present more details in the Appendix to describe the complete supergravity action, including scalars, fermions, and vectors. In supergravity in Kallosh:2015sea and here in eq. (2) in absence of fermions
@@ 3 Consistency issue of streamlined supergravity and its resolution
$K(T,\bar{T},s,\bar{s})=-3\alpha\ln(T+\bar{T})+\frac{F_{s}\bar{F}_{\bar{s}}}{(T+\bar{T})^{3\alpha}V(T,\bar{T})+3|W_{0}|^{2}(1-\alpha)}\,s\,\bar{s}\ ,$ (14)
@@
${{\cal L}(T,\bar{T})\over\sqrt{-g}}={R\over 2}-{3\alpha\over 4}\,{\partial T\partial\bar{T}\over({\rm Re}\,T)^{2}}-V(T,\bar{T})\ .$ (15)
@@
However, we can use local supersymmetry to gauge-fix the fermion from the nilpotent multiplet to vanish,
@@
Therefore, the kinetic terms of the boson $s$ and the fermion $\psi^{s}$ with a negative $K_{s\bar{s}}$ are not harmful. There is no vacuum instability due to these fermions since their kinetic term is either absent in $\psi^{s}=0$ gauge, or involves terms with negative $K_{s\bar{s}}$ which are not quadratic in physical fermions since they contain coupling to moduli $z^{i}$. Thus the streamlined supergravity defined by $K_{s\bar{s}}$ in eqs. (11) is consistent even in the models where $K_{s\bar{s}}<0$ at some points in moduli space.
@@ 4.1.1 E-models
For models with an exponential approach to the plateau, consider the case of E-models of $\alpha$-attractors Ferrara:2013rsa ; Kallosh:2013yoa . In streamlined supergravity, these models are described by complex half-plane variables $T=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$ in hyperbolic geometry with the nilpotent superfield $S$ with $S^{2}=0$. The simplest models of this class have potentials with a plateau, which is approached exponentially fast at large $\varphi$. In particular, one may consider models (14), (15) with
$V_{E}(T,\bar{T})=\Lambda+V_{0}(1-T)^{n}(1-\bar{T})^{n}=\Lambda+V_{0}\Big(1-e^{-\sqrt{2\over 3\alpha}\varphi}\Big)^{2n}\ .$ (24)
Here $T=\bar{T}=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$, and inflation occurs at large values of the canonically normalized field $\varphi$. To stabilize the axion field at $a=0$, one may simply add to this potential a term $m^{2}(T,\bar{T})(T-\bar{T})^{2}$, which vanishes for $T=\bar{T}$. (Note also that in this class of theories, one can construct consistent $\alpha$-attractor models with the same cosmological predictions without stabilizing the axion potential Achucarro:2017ing ; Linde:2018hmx ).
These models have a potential with a plateau, which is approached exponentially fast at large $\varphi$.
$V_{E}(T,\bar{T})=V_{0}\Big(1-2n\,e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)\ .$ (25)
The factor 2n in the last equation can be absorbed by a redefinition (a shift) of the field $\varphi$. This model can also be interpreted as pole inflation Galante:2014ifa ; Terada:2016nqg , with the kinetic terms having a pole of second order at $T=\bar{T}\to 0$ (15).
@@ 4.1.2 T-models
The streamlined supergravity version of $\alpha$-attractors in disk variables, based on the general equation (4), corresponds to
$K(Z,\bar{Z},s,\bar{s})=$ $-3\alpha\ln(1-Z\bar{Z})+\frac{F_{s}\bar{F}_{\bar{s}}}{(1-Z\bar{Z})^{3\alpha}\,V(Z,\bar{Z})+3|W_{0}|^{2}(1-\alpha Z\bar{Z})}s\bar{s}\ .$ (28)
@@
We assume that $\Lambda\ll V_{0}$, so one can ignore $\Lambda$ during inflation. The potential in disk variables with the stabilized axion and $Z=\bar{Z}=\tanh{\varphi\over\sqrt{6\alpha}}$ during inflation at large $\varphi$ is
$V(Z,\bar{Z})=V_{0}\tanh^{2n}{\varphi\over\sqrt{6\alpha}}=V_{0}\Big(1-2n\,e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)\ .$ (32)
We find the same exponential approach to the plateau in terms of a canonical field $\varphi$ as in the E-model (25). The factor 2n in the last equation can be absorbed by a redefinition (a shift) of the field $\varphi$. This is a case of the pole inflation in disk variables, where the kinetic terms have a pole of the order $p=2$ at the boundary $Z\bar{Z}\to 1$.
The supergravity version of polynomial $\alpha$-attractors Kallosh:2022feu is also described in (14), (15), with the kinetic term with the pole of order $p=2$. The potential in this case has particular features near the pole, where, e.g., for $k=2,4$
$V_{\rm pole}^{\rm ln}(T,\bar{T})\to V_{0}\Big(1-\Big({2\over\ln T\bar{T}}\Big)^{k}+\dots\Big)=V_{0}\Big(1-\Big({\sqrt{2\over 3\alpha}\varphi}\Big)^{-k}+\dots\Big)\ .$ (33)
In this model, the plateau at large $\varphi$ is approached polynomially Kallosh:2022feu ; Kallosh:2025ijd .
[…]
4.2.1 Pole inflation in half-plane variables
We now describe pole inflation with kinetic terms that have a pole of order $p>2$. To do it, we will take the following function $K(T,\bar{T},s,\bar{s})$ defined in the general equation (4):
$K(T,\bar{T},s,\bar{s})={\mu^{2}\over 2}({T}\overline{T})^{2-p\over 2}+{F_{s}\bar{F}_{\bar{s}}\over{V}\left({T},\overline{{T}}\right)e^{-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}}+W_{0}^{2}\left(3-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}\right)}\,s\bar{s}$ (34)
@@ 4.2.1 Pole inflation in half-plane variables
This model describes pole inflation Galante:2014ifa ; Terada:2016nqg with $p>2$, where in the canonical variables
$V|_{\varphi\to\infty}\to V_{0}(1-\Big({\mu\over\varphi}\Big)^{k}+\dots)\ ,\qquad k={2\over p-2}>0\ .$ (37)
4.2.2 Pole inflation in disk variables
The streamlined supergravity version of pole inflation with pole order $p>2$ in disk variables can be described by
$K(Z,\bar{Z},s,\bar{s})=$ ${\mu^{2}\over 2}(1-Z\bar{Z})^{2-p}+\frac{F_{s}\bar{F}_{\bar{s}}}{e^{{\mu^{2}\over 2}(1-Z\bar{Z})^{2-p}}\,V(Z,\bar{Z})+|W_{0}|^{2}\Big(3-{\mu^{2}\over 2}{(p-2)Z\bar{Z}(1-Z\bar{Z})^{p-2}\over 1+(p-2)Z\bar{Z}}\Big)}s\bar{s}\ .$ (38)
${{\cal L}(Z,\bar{Z})\over\sqrt{-g}}={R\over 2}-{A_{p}\over 2}\,{\partial Z\partial\bar{Z}\over(1-Z\bar{Z})^{p}}-V(Z,\bar{Z})\ ,\qquad A_{p}={\mu^{2}}(p-2)\big(1+(p-2)Z\bar{Z}\big)\ .$ (39)
This is the pole-inflation model in which the kinetic terms have a pole of order $p>2$ at the boundary $Z\bar{Z}\to 1$. The polynomial approach to the plateau in terms of the canonical field $\varphi$ is similar to the one in the model (34). Observational results of these inflationary models depend only on $p$, on the value of the potential $V(Z,\bar{Z})$ at $Z\bar{Z}\to 1$, and on the value of $A_{p}$ at $Z\bar{Z}\to 1$.
$a_{p}=A_{p}\Big|_{Z\bar{Z}\to 1}=\mu^{2}(p-2)(p-1)\ .$ (40)
@@ 5 Discussion
In this paper we presented an efficient and relatively simple version of the streamlined $N=1$ supergravity with chiral multiplets and a nilpotent multiplet $S^{2}=0$ with the general type Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ leading to any desired scalar potential $V(z^{i},\bar{z}^{\bar{i}})$, once the proper choice of the Kähler metric of the nilpotent multiplet is made. The total Kähler potential, superpotential, and the bosonic action of this new supergravity are given in eqs. (12), (13). We provided some examples illustrating our methods in applications to inflationary potentials.
@@ Acknowledgement
We acknowledge many stimulating discussions over the years on supergravity with nilpotent multiplets with I. Antoniadis, J.J. Carrasco, S. Ferrara, D. Freedman, E. McDonough, D. Roest, M. Scalisi, A. Van Proeyen, T. Wrase, and Y. Yamada. This work was supported by the Leinweber Institute for Theoretical Physics at Stanford and by NSF Grant PHY-2310429.